# Universal logical operations with a dynamical qubit in Floquet code

### v1

@@ III Discussion and outlook
In this work, we implement the FBS code on a superconducting quantum processor, encoding a dynamical logical qubit in addition to a conventional static logical qubit. We experimentally demonstrate universal two-qubit logical operations integrated with repetitive quantum error detection, including encoding, measurement, arbitrary single-qubit rotations, and a CNOT gate. Some of these logical operations are implemented in a FT manner and outperform their nFT counterparts. Our first demonstration of universal logical operations with a dynamical logical qubit highlights the significant role played by Floquet codes in achieving resource-efficient FT quantum computation.
There are several challenges that remain for future work: (1) Despite improvements in code rate using the FBS code, our experiments show that the dynamical logical qubit is more susceptible to errors, requiring a larger code distance for effective protection. Specifically, the code distance for the dynamical qubit is approximately $d/2$ when the FBS code is implemented on a $d\times d$ square lattice [42]. Further research is needed to scale up the FBS code to larger lattices to enable effective error correction for the dynamical qubit. (2) Implementing FT entanglement and non-Clifford gates with the dynamical logical qubit is a critical next step. It has been shown that standard methods used in surface code, such as lattice surgery [24] and braiding twist defects [38], can be adapted to Floquet codes. A recent theoretical work has explored FT logical operations using a generalized Floquet code in a honeycomb lattice, demonstrating that entanglement operations can be achieved by simply reordering the measurement schedule [66]. This suggests that Floquet codes may offer a simpler and more natural pathway to FT quantum computing.
@@ IV.1 Logical operator definition in the FBS code
For Pauli operators $P\in\{X,Y,Z\}$, the sign at round $r$ is given by formula $\Gamma_{P}^{(r)}=\prod_{i=1}^{r}\gamma_{P}^{(i)}$, where $\gamma_{P}^{(i)}$ is defined in Table 2. Combining all of the above, the logical $P$ operator of the dynamical qubit at round $r$ is given by $P_{\rm d}^{(r)}=\Gamma_{P}^{(r)}P_{\rm d}^{\rm Q},$ where ${\rm Q}={\rm A},{\rm B},{\rm C},{\rm D}$ for $r\equiv 1,2,3,0\pmod{4}$, respectively.
@@ IV.3 Two-qubit Logical Quantum State Tomography
We optimize the logical density matrix $\rho_{\rm L}$ using the convex optimization package cvxpy [67], constraining the density matrix to be Hermitian, unit-trace and positive semi-definite. The fidelity of the reconstruted state relative to a target pure state is calculated as ${\rm Tr}(\rho_{\rm L}\rho_{\rm ideal})$.
@@ IV.4 Two-qubit Logical Quantum Process tomography
The LPTM $R$, a $16\times 16$ matrix, maps an input state vector $\vec{p}$ to an output state vector $\vec{p^{\prime}}=R\vec{p}$ [68]. To reconstruct the LPTM, we employ convex optimization to find an optimal LPTM that satisfies all the constraints of a physical quantum channel, using an overcomplete set of input states.
@@ Acknowledgements
We thank Bujiao Wu for valuable discussions. This work was supported by the Innovation Program for Quantum Science and Technology (2021ZD0301703), the Science, Technology and Innovation Commission of Shenzhen Municipality (KQTD20210811090049034), the National Natural Science Foundation of China (12174178, 12374474, 123b2071).

### v2

@@ III Discussion and outlook
In this work, we implement the FBS code on a superconducting quantum processor, encoding a dynamical logical qubit in addition to a conventional static logical qubit. We experimentally demonstrate universal two-qubit logical operations integrated with repetitive quantum error detection, including encoding, measurement, arbitrary single-qubit rotations, and a CNOT gate. Some of these logical operations are implemented in a FT manner. Our first demonstration of universal logical operations with a dynamical logical qubit highlights the significant role played by Floquet codes in achieving resource-efficient FT quantum computation.
There are several challenges that remain for future work: (1) Despite improvements in code rate using the FBS code, our experiments show that the dynamical logical qubit is more susceptible to errors, requiring a larger code distance for effective protection. Specifically, the code distance for the dynamical qubit is approximately $d/2$ when the FBS code is implemented on a $d\times d$ square lattice [44]. Further research is needed to scale up the FBS code to larger lattices to enable effective error correction for the dynamical qubit. (2) Implementing FT entanglement and non-Clifford gates with the dynamical logical qubit is a critical next step. It has been shown that standard methods used in surface code, such as lattice surgery and braiding twist defects, can be adapted to Floquet codes [25, 40]. A recent theoretical work has explored FT logical operations using a generalized Floquet code in a honeycomb lattice, demonstrating that entanglement operations can be achieved by simply reordering the measurement schedule [73]. This suggests that Floquet codes may offer a simpler and more natural pathway to FT quantum computing.
@@ IV.1 Logical operator definition in the FBS code
For Pauli operator $P\in\{X,Y,Z\}$, its sign at round $r$ is given by formula $\Gamma_{P}^{(r)}=\prod_{i=1}^{r}\gamma_{P}^{(i)}$, where $\gamma_{P}^{(i)}$ is defined in Table 2. Combining all of the above, the logical $P$ operator of the dynamical qubit at round $r$ is given by $P_{\rm d}^{(r)}=\Gamma_{P}^{(r)}P_{\rm d}^{\rm Q},$ where ${\rm Q}={\rm A},{\rm B},{\rm C},{\rm D}$ for $r\equiv 1,2,3,0\pmod{4}$, respectively.
@@ IV.3 Two-qubit Logical Quantum State Tomography
We optimize the logical density matrix $\rho_{\rm L}$ using the convex optimization package cvxpy [74], constraining the density matrix to be Hermitian, unit-trace and positive semi-definite. The fidelity of the reconstructed state relative to a target pure state is calculated as ${\rm Tr}(\rho_{\rm L}\rho_{\rm ideal})$.
@@ IV.4 Two-qubit Logical Quantum Process tomography
The LPTM $R$, a $16\times 16$ matrix, maps an input state vector $\vec{p}$ to an output state vector $\vec{p^{\prime}}=R\vec{p}$ [72]. We reconstruct the LPTM using a complete set of input-output state vectors. Convex optimization is employed to find an optimal LPTM that satisfies all the constraints of a physical quantum channel.
@@ Acknowledgements
We thank Bujiao Wu for valuable discussion. This work was supported by the Innovation Program for Quantum Science and Technology (2021ZD0301703), the Science, Technology and Innovation Commission of Shenzhen Municipality (KQTD20210811090049034), the National Natural Science Foundation of China (12174178, 12374474, 123b2071).